# Teaching Large Language Models an Unseen Language on the Fly

### v1

@@ Appendix D Example Output
Appendix E Experiment on MTOB
@@
Appendix F Learning Syntax Explicitly
@@
We conduct a pilot study to investigate whether LLMs can comprehend the the syntactical information explicitly given in the prompt. We focus on the order of modifiers and modified elements, which differs greatly between Zhuang and Chinese. We select 10 zh2za testing samples where GPT-3.5-Turbo fails on the order of modifiers and modified elements, and try different strategies to incorporate this syntactical rule into the prompt.
@@
First, we attempt to directly declare a rule of word order between modifiers and modified elements in the prompt, i.e., different from Chinese, Zhuang puts the modifiers after the modified elements. GPT-3.5-Turbo correctly addresses only 1 of the 10 testing examples.
@@
Second, we use chain-of-thought (CoT) reasoning to encourage the model to analyze the word order in the given source sentence. The format of CoT is shown in Table 13. We notice that GPT-3.5-Turbo often fails to identify the modifiers and the modified elements from the given sentence, probably because of a lack of general linguistic knowledge. After we hint the modifiers and the modified elements of the given sentences in the prompt, GPT-3.5-Turbo can correctly analyze and address their order for 7 out of 10 testing examples. This shows the potential of explicitly teaching LLMs syntactic rules through CoT.
Appendix G User Study
@@
Data. From the easy subset of ZhuangBench, we sample 20 instances for zh2za and 40 instances for za2zh. It is more difficult to perform translations from high- to low-resource languages, so we use fewer zh2za instances than za2zh ones to reduce the burden on the participants.
@@
Interface. We develop an interface for the user study, as shown in Figure 2. To reduce the participants’ burden of dictionary searching in the Human Only setting, we also show the prompt generated by DiPMT++ to our participants for reference, which already contains possible translations for each word in the sentence to be translated. The time used for translation and the actions of the users are automatically tracked. During the study, the translation instances with and without LLM assistance appear alternately.
@@
Participants. Before the study, we inform the participants of the scientific use of their translation data and obtain consent from them. The participants are adequately paid given their demographic.
@@
Human Actions. In Table 14, we show an example of how the participant switches between searching Chinese and Zhuang words/phrases during zh2za translation.
@@
Chinese: 被人骗了,损失了100多万元。 (I was deceived, and lost more than 1 million yuan.)
@@
Chinese: 从公共场所回来、捂着嘴巴咳嗽、吃饭之前,我们都要洗手,防止患病。(When we come back from public places, cough with our mouths covered, and before eating, we all wash our hands to prevent getting sick.)
Zhuang: Daj giz vunz lai dauqma、 goemq bak ae、 yaek gwn haeux, raeuz cungj aeu swiq fwngz, fuengzre baenzbingh.
@@
Chinese: 我先是住在监狱旁边一个客店里的,初冬已经颇冷,蚊子却还多,后来用被盖了全身,用衣服包了头脸,只留两个鼻孔出气。 (I first lived in an inn next to the prison, and it was quite cold in the early winter, but there were still many mosquitoes, and then I covered my whole body with a quilt, covered my head and face with clothes, and left only two nostrils to breathe.)
@@
# 请仿照样例,参考给出的词汇,将汉语句子翻译成壮语。 (Please follow the example and refer to the given vocabulary to translate the Chinese sentences into Zhuang.)
## 请将下面的汉语句子翻译成壮语:好。明天你就要回去了,今天晚上我让我妻子弄几个菜,咱们喝两杯。(Please translate the following Chinese sentence into Zhuang: OK. You’re going back tomorrow. I’ll ask my wife to prepare some dishes tonight and we’ll have a drink.)
## 在上面的句子中,汉语词语“好”在壮语对应的词是“ndei”或“baenz”;汉语词语“明天”在壮语对应的词是“ngoenzcog”或“ngoenzbyug”;… (In the above sentence, the Chinese word “good” corresponds to the Zhuang word “ndei” or“baenz”; the Chinese word “tomorrow” corresponds to the Zhuang word “ngoenzcog” or “ngoenzbyug”;…)
## 所以,完整的壮语翻译是:Ndei. Ngoenzcog mwngz couh yaek baema lo, haemhneix gou heuh yah gou loengh geij yiengh byaek, raeuz ndoet song cenj. (So, the complete Zhuang translation is: Ndei. Ngoenzcog mwngz couh yaek baema lo, haemhneix gou heuh yah gou loengh geij yiengh byaek, raeuz ndoet song cenj.)
@@
## 请将下面的汉语句子翻译成壮语:现在,农村许多老年人年轻人会使用手机来买卖商品,很方便。(Please translate the following Chinese sentence into Zhuang: Nowadays, many elderly and young people in rural areas use mobile phones to buy and sell goods, which is very convenient.)
## 在上面的句子中,汉语词语“现在”在壮语对应的词是“seizneix”或“neix”;… (In the above sentence, the Chinese word “nowadays” corresponds to the Zhuang word “seizneix” or“neix”; …)
## 所以,完整的壮语翻译是: (So, the complete Zhuang translation is:)
@@
Input:
Seizneix, lajmbanj haujlai bouxlaux bouxcoz rox yungh soujgih daeuj gaicawx doxgaiq, fuengbienh lai lo.
Gold:
现在,农村许多老年人年轻人会使用手机来买卖商品,很方便。(Nowadays, many elderly and young people in rural areas use mobile phones to buy and sell goods, which is very convenient.)
mT5-large (Finetune):
现在,农村贫困人口多,贫困人口多,贫困人口多。 (Nowadays, there are many poor people in rural areas, there are many poor people, there are many poor people.)
LLaMA-2-7B-chat (Finetune):
现在,农村多少少的农车轮到市场上,便利了。 (Nowadays, more and more farm wheels from rural areas have come to the market, which is convenient.)
LLaMA-2-7B-chat (DiPMT++):
现在我们可以通过手机来买卖,方便很多。 (Now we can buy and sell through mobile phones, which is much more convenient.)
Qwen-14B-chat (DiPMT):
现在,乡下的父老、青年都知道用东西买卖很方便。 (Nowadays, the elders and young people in the countryside know that it is very convenient to buy and sell things.)
Qwen-14B-chat (DiPMT++):
现在,乡下的老人和青年都知道用手机买卖东西很方便。 (Nowadays, old people and young people in the countryside know that it is convenient to buy and sell things using mobile phones.)
GPT-4 (Direct):
自从,我们村子里来了许多外来人口,村里的面貌变化很大,蓬勃发展起来了。 (Since then, many outsiders have come to our village. The appearance of the village has changed a lot and it has developed vigorously.)
GPT-4 (DiPMT++):
现在,农村里许多老人和青年都知道用手机买卖东西了,方便多了。 (Nowadays, many old people and young people in rural areas know how to use mobile phones to buy and sell things, which is much more convenient.)
@@
Input:
现在,农村许多老年人年轻人会使用手机来买卖商品,很方便。 (Nowadays, many elderly and young people in rural areas use mobile phones to buy and sell goods, which is very convenient.)
Gold:
Seizneix, lajmbanj haujlai bouxlaux bouxcoz rox yungh soujgih daeuj gaicawx doxgaiq, fuengbienh lai lo.
mT5-large (Finetune):
Aen gvaq, aen gvaq ndawbiengz gunghyez gunghyez cungj aeu ndawbiengz cungj aeu yungh ndawbiengz cungj aeu yungh ndawbiengz ndawbiengz daeuj ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej nda
LLaMA-2-7B-chat (Finetune):
Seizneix, lajmbanj haujlai bouxgeq ndaej sawjyungh soujgih daeuj cawx gaiq sang, ndaej fuengbienh.
LLaMA-2-7B-chat (DiPMT++):
Seizneix hauqgai sanghbinj gaicawx maz daeuj, lai fuengbienh.
Qwen-14B-chat (DiPMT):
Seizneix, lajmbanj haujlai neix, caetneix年轻人 lox sawjyungh caek maz gaicawx, lai fuengbienh.
Qwen-14B-chat (DiPMT++):
Seizneix lajmbanj haujlai bouxgeq mbanjveiz rox yungh soujgih daeuj gaicawx sanghbinj lai fuengbienh.
GPT-4 (Direct):
Yienvaiz, pwngzloengh bouxlaeuz sw’ciengz niengh yinzhangz daenghhoz gwn lajmbouj raen raemmbawz, ndei ndaw.
GPT-4 (DiPMT++):
Seizneix, lajmbanj miz haujlai bouxgeq mbanjveiz rox yungh soujgih daeuj gaicawx sanghbinj, lai bienh.
@@
## 请将下面的汉语句子翻译成壮语(翻译时请注意,与汉语不同,壮语中的修饰语放在被修饰的中心词之后):我们村有一条河。(Please translate the following Chinese sentence into Zhuang (Please note when translating that, unlike Chinese, the modifier in Zhuang language is placed after the modified elements): There is a river in our village.)
## 在上面的句子中,汉语词语“我们”在壮语对应的词是“dou”;汉语词语“村”在壮语对应的词是“mbanj”;… (In the above sentence, the Chinese word “our” corresponds to the Zhuang word “dou”; the Chinese word “village” corresponds to the Zhuang word “mbanj”;…)
## 那么让我们来一步一步的翻译: (So let’s translate it step by step:)
在该句中,存在修饰语和被修饰语,修饰语是“我们”,被修饰语是“村”,“我们”翻译成“gou”,“村”翻译成“mbanj”。 (In this sentence, there are modifiers and modified elements. The modifier is “our” and the modified element is “village”. “Our” is translated into "gou" and “village” is translated into “mbanj”.)
在壮语中,修饰语放在被修饰的中心词之后,所以“我们村”翻译成“mbaij dou”。 (In Zhuang, the modifier is placed after the modified element, so "our village" is translated as "mbaij dou".)
基于上述分析,该汉语句子的最终壮语翻译是:Mbanj dou miz diuz dah. (Based on the above analysis, the final Zhuang translation of this Chinese sentence is: Mbanj dou miz diuz dah.)
@@
Source Sentence: 这个果园种有多少种果树?(How many types of fruit trees are planted in this orchard?)
@@
Word Search
Chinese
种类
Word Search
Chinese
Corpus Search
Chinese
多少种
Corpus Search
Chinese
这个果园
Corpus Search
Chinese
这个
Word Search
Zhuang
aen
Word Search
Zhuang
neix
Word Search
Zhuang
ndaem
Word Search
Zhuang
cungj

### v2

@@ Appendix D Example Output
Appendix E Experiments on MTOB
@@
Appendix F Experiments on Seen Languages
We also evaluate DiPMT++ on more low-resource languages that might have been seen during pre-training, i.e., the languages that LLMs can translate with a non-zero BLEU score by zero-shot prompting. In this way, we could show a clear picture of the applicability of our method.
Languages
We use 7 low-resource languages including Estonian (et), Lithuanian (lt), Latvian (lv), Macedonian (mk), Slovak (sk), Albanian (sq), and Filipino (tl).
Setup
For each language, we use the 2,009 publicly available sentences from Flores-200 NLLB Team et al. (2022). We sample 200 sentences from it for testing and the rest form the corpus for retrieval. We use the dictionaries from MUSE Lample et al. (2018), consisting of 5K entries. We use Llama-2-13B-chat as the backbone model.
Results
We report the results in Table 9. The backbone model already has varied abilities regarding the studied languages, achieving non-zero performance through zero-shot direct prompting. This confirms that the model might already see texts in these languages during pretraining. Regarding the very low-resource languages, of which the model’s ability is weak (BLEU < 10), such as Estonian, Lithuanian, Latvian, and Albanian, DiPMT++ outperforms DiPMT and direct prompting by a large margin. In terms of the mid-resource languages, of which the model has a decent ability (BLEU > 20), such as Afrikaans, Indonesian, and Malay, the advantage of DiPMT++ is less pronounced. In conclusion, our method also works on low-resource languages already seen by LLMs, especially on the extremely low-resource ones, of which the model’s original ability is poor.
Method et lt lv mk sk sq tl
en2et et2en en2lt lt2en en2lv lv2en en2mk mk2en en2sk sk2en en2sq sq2en en2tl tl2en
Direct 5.3 12.0 4.5 10.6 4.2 10.4 11.3 31.4 12.2 32.5 3.5 9.8 13.5 27.8
DiPMT 8.1 16.6 6.7 15.5 7.0 17.0 15.0 34.2 14.4 34.7 5.8 15.9 16.9 32.6
DiPMT++ 9.1 18.3 7.7 17.0 8.8 18.2 14.6 34.2 14.4 35.4 9.7 17.5 18.3 31.5
Table 9: Performance of different methods on languages seen by Llama-2-13B-chat. We use 3-shot exemplars for DiPMT and DiPMT++. The best scores are made bold.
Appendix G Learning Syntax Explicitly
@@
We conduct a pilot study to investigate whether LLMs can comprehend the the syntactical information explicitly given in the prompt. We focus on the order of modifiers and modified elements, which differs greatly between Zhuang and Chinese. We select 10 zh2za testing samples where GPT-3.5 fails on the order of modifiers and modified elements, and try different strategies to incorporate this syntactical rule into the prompt.
@@
First, we attempt to directly declare a rule of word order between modifiers and modified elements in the prompt, i.e., different from Chinese, Zhuang puts the modifiers after the modified elements. GPT-3.5 correctly addresses only 1 of the 10 testing examples.
@@
Second, we use chain-of-thought (CoT) reasoning to encourage the model to analyze the word order in the given source sentence. The format of CoT is shown in Table 14. We notice that GPT-3.5 often fails to identify the modifiers and the modified elements from the given sentence, probably because of a lack of general linguistic knowledge. After we hint the modifiers and the modified elements of the given sentences in the prompt, GPT-3.5 can correctly analyze and address their order for 7 out of 10 testing examples. This shows the potential of explicitly teaching LLMs syntactic rules through CoT.
Appendix H User Study
Data
@@ Appendix H User Study
From the easy subset of ZhuangBench, we sample 20 instances for zh2za and 40 instances for za2zh. It is more difficult to perform translations from high- to low-resource languages, so we use fewer zh2za instances than za2zh ones to reduce the burden on the participants.
Interface
@@
We develop an interface for the user study, as shown in Figure 2. To reduce the participants’ burden of dictionary searching in the Human Only setting, we also show the prompt generated by DiPMT++ to our participants for reference, which already contains possible translations for each word in the sentence to be translated. The time used for translation and the actions of the users are automatically tracked. During the study, the translation instances with and without LLM assistance appear alternately.
Participants
@@
Before the study, we inform the participants of the scientific use of their translation data and obtain consent from them. The participants are adequately paid given their demographic.
Human Actions
@@
In Table 15, we show an example of how the participant switches between searching Chinese and Zhuang words/phrases during zh2za translation.
Performance Change During the Study
We do not observe significant performance changes as the participants translate more sentences, since each participant is only assigned 20 instances during the task. Yet, we find that the participants spend slightly more time on the first few instances and their translation speed becomes steady afterwards.
Based on our study, we observe that the benefit of having LLMs utility persists through the translation process. Although humans might become more familiar with the interface, after conducting more translations, it is unlikely for them to accurately remember all the lexical knowledge embedded in these instances, let alone complex syntactical structures. Therefore, LLMs can consistently assist in providing suggestions related to word choice, collocations, and local syntactical structures.
@@
{CJK*}UTF8gbsnChinese: 被人骗了,损失了100多万元。 (I was deceived, and lost more than 1 million yuan.)
@@
{CJK*}UTF8gbsnChinese: 从公共场所回来、捂着嘴巴咳嗽、吃饭之前,我们都要洗手,防止患病。(When we come back from public places, cough with our mouths covered, and before eating, we all wash our hands to prevent getting sick.)
Zhuang: Daj giz vunz lai dauqma{CJK*}UTF8gbsn、 goemq bak ae{CJK*}UTF8gbsn、 yaek gwn haeux, raeuz cungj aeu swiq fwngz, fuengzre baenzbingh.
@@
{CJK*}UTF8gbsnChinese: 我先是住在监狱旁边一个客店里的,初冬已经颇冷,蚊子却还多,后来用被盖了全身,用衣服包了头脸,只留两个鼻孔出气。 (I first lived in an inn next to the prison, and it was quite cold in the early winter, but there were still many mosquitoes, and then I covered my whole body with a quilt, covered my head and face with clothes, and left only two nostrils to breathe.)
@@
{CJK*}UTF8gbsn# 请仿照样例,参考给出的词汇,将汉语句子翻译成壮语。 (Please follow the example and refer to the given vocabulary to translate the Chinese sentences into Zhuang.)
{CJK*}UTF8gbsn## 请将下面的汉语句子翻译成壮语:好。明天你就要回去了,今天晚上我让我妻子弄几个菜,咱们喝两杯。(Please translate the following Chinese sentence into Zhuang: OK. You’re going back tomorrow. I’ll ask my wife to prepare some dishes tonight and we’ll have a drink.)
{CJK*}UTF8gbsn## 在上面的句子中,汉语词语“好”在壮语对应的词是“ndei”或“baenz”;汉语词语“明天”在壮语对应的词是“ngoenzcog”或“ngoenzbyug”;… (In the above sentence, the Chinese word “good” corresponds to the Zhuang word “ndei” or“baenz”; the Chinese word “tomorrow” corresponds to the Zhuang word “ngoenzcog” or “ngoenzbyug”;…)
{CJK*}UTF8gbsn## 所以,完整的壮语翻译是:Ndei. Ngoenzcog mwngz couh yaek baema lo, haemhneix gou heuh yah gou loengh geij yiengh byaek, raeuz ndoet song cenj. (So, the complete Zhuang translation is: Ndei. Ngoenzcog mwngz couh yaek baema lo, haemhneix gou heuh yah gou loengh geij yiengh byaek, raeuz ndoet song cenj.)
@@
{CJK*}UTF8gbsn## 请将下面的汉语句子翻译成壮语:现在,农村许多老年人年轻人会使用手机来买卖商品,很方便。(Please translate the following Chinese sentence into Zhuang: Nowadays, many elderly and young people in rural areas use mobile phones to buy and sell goods, which is very convenient.)
{CJK*}UTF8gbsn## 在上面的句子中,汉语词语“现在”在壮语对应的词是“seizneix”或“neix”;… (In the above sentence, the Chinese word “nowadays” corresponds to the Zhuang word “seizneix” or“neix”; …)
{CJK*}UTF8gbsn## 所以,完整的壮语翻译是: (So, the complete Zhuang translation is:)
@@
Input: Seizneix, lajmbanj haujlai bouxlaux bouxcoz rox yungh soujgih daeuj gaicawx doxgaiq, fuengbienh lai lo.
Gold: {CJK*}UTF8gbsn现在,农村许多老年人年轻人会使用手机来买卖商品,很方便。(Nowadays, many elderly and young people in rural areas use mobile phones to buy and sell goods, which is very convenient.)
mT5-large (Finetune): {CJK*}UTF8gbsn现在,农村贫困人口多,贫困人口多,贫困人口多。 (Nowadays, there are many poor people in rural areas, there are many poor people, there are many poor people.)
Llama-2-7B-chat (Finetune): {CJK*}UTF8gbsn现在,农村多少少的农车轮到市场上,便利了。 (Nowadays, more and more farm wheels from rural areas have come to the market, which is convenient.)
Llama-2-7B-chat (DiPMT++): {CJK*}UTF8gbsn现在我们可以通过手机来买卖,方便很多。 (Now we can buy and sell through mobile phones, which is much more convenient.)
Qwen-14B-chat (DiPMT): {CJK*}UTF8gbsn现在,乡下的父老、青年都知道用东西买卖很方便。 (Nowadays, the elders and young people in the countryside know that it is very convenient to buy and sell things.)
Qwen-14B-chat (DiPMT++): {CJK*}UTF8gbsn现在,乡下的老人和青年都知道用手机买卖东西很方便。 (Nowadays, old people and young people in the countryside know that it is convenient to buy and sell things using mobile phones.)
GPT-4 (Direct): {CJK*}UTF8gbsn自从,我们村子里来了许多外来人口,村里的面貌变化很大,蓬勃发展起来了。 (Since then, many outsiders have come to our village. The appearance of the village has changed a lot and it has developed vigorously.)
GPT-4 (DiPMT++): {CJK*}UTF8gbsn现在,农村里许多老人和青年都知道用手机买卖东西了,方便多了。 (Nowadays, many old people and young people in rural areas know how to use mobile phones to buy and sell things, which is much more convenient.)
@@
Input: {CJK*}UTF8gbsn现在,农村许多老年人年轻人会使用手机来买卖商品,很方便。 (Nowadays, many elderly and young people in rural areas use mobile phones to buy and sell goods, which is very convenient.)
Gold: Seizneix, lajmbanj haujlai bouxlaux bouxcoz rox yungh soujgih daeuj gaicawx doxgaiq, fuengbienh lai lo.
mT5-large (Finetune): Aen gvaq, aen gvaq ndawbiengz gunghyez gunghyez cungj aeu ndawbiengz cungj aeu yungh ndawbiengz cungj aeu yungh ndawbiengz ndawbiengz daeuj ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej ndaej nda
Llama-2-7B-chat (Finetune): Seizneix, lajmbanj haujlai bouxgeq ndaej sawjyungh soujgih daeuj cawx gaiq sang, ndaej fuengbienh.
Llama-2-7B-chat (DiPMT++): Seizneix hauqgai sanghbinj gaicawx maz daeuj, lai fuengbienh.
Qwen-14B-chat (DiPMT): Seizneix, lajmbanj haujlai neix, caetneix{CJK*}UTF8gbsn年轻人 lox sawjyungh caek maz gaicawx, lai fuengbienh.
Qwen-14B-chat (DiPMT++): Seizneix lajmbanj haujlai bouxgeq mbanjveiz rox yungh soujgih daeuj gaicawx sanghbinj lai fuengbienh.
GPT-4 (Direct): Yienvaiz, pwngzloengh bouxlaeuz sw’ciengz niengh yinzhangz daenghhoz gwn lajmbouj raen raemmbawz, ndei ndaw.
GPT-4 (DiPMT++): Seizneix, lajmbanj miz haujlai bouxgeq mbanjveiz rox yungh soujgih daeuj gaicawx sanghbinj, lai bienh.
@@
{CJK*}UTF8gbsn## 请将下面的汉语句子翻译成壮语(翻译时请注意,与汉语不同,壮语中的修饰语放在被修饰的中心词之后):我们村有一条河。(Please translate the following Chinese sentence into Zhuang (Please note when translating that, unlike Chinese, the modifier in Zhuang language is placed after the modified elements): There is a river in our village.)
{CJK*} UTF8gbsn## 在上面的句子中,汉语词语“我们”在壮语对应的词是“dou”;汉语词语“村”在壮语对应的词是“mbanj”;… (In the above sentence, the Chinese word “our” corresponds to the Zhuang word “dou”; the Chinese word “village” corresponds to the Zhuang word “mbanj”;…)
{CJK*} UTF8gbsn## 那么让我们来一步一步的翻译: (So let’s translate it step by step:)
{CJK*}UTF8gbsn在该句中,存在修饰语和被修饰语,修饰语是“我们”,被修饰语是“村”,“我们”翻译成“gou”,“村”翻译成“mbanj”。 (In this sentence, there are modifiers and modified elements. The modifier is “our” and the modified element is “village”. “Our” is translated into "gou" and “village” is translated into “mbanj”.)
{CJK*}UTF8gbsn在壮语中,修饰语放在被修饰的中心词之后,所以“我们村”翻译成“mbaij dou”。 (In Zhuang, the modifier is placed after the modified element, so "our village" is translated as "mbaij dou".)
{CJK*}UTF8gbsn基于上述分析,该汉语句子的最终壮语翻译是:Mbanj dou miz diuz dah. (Based on the above analysis, the final Zhuang translation of this Chinese sentence is: Mbanj dou miz diuz dah.)
@@
Source Sentence: {CJK*}UTF8gbsn这个果园种有多少种果树?(How many types of fruit trees are planted in this orchard?)
@@
Word Search Chinese {CJK*}UTF8gbsn种类
Word Search Chinese {CJK*}UTF8gbsn种
Corpus Search Chinese {CJK*}UTF8gbsn多少种
Corpus Search Chinese {CJK*}UTF8gbsn这个果园
Corpus Search Chinese {CJK*}UTF8gbsn这个
Word Search Zhuang aen
Word Search Zhuang neix
Word Search Zhuang ndaem
Word Search Zhuang cungj